OpenStudy (anonymous):

can someone help me solve for x. 3x^2=-7x-2

OpenStudy (anonymous):

Do you know Bhaskara?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

x = -1/3

OpenStudy (anonymous):

\[ x =- \frac{1}{3} \]

OpenStudy (anonymous):

had lunch with him yesterday

OpenStudy (anonymous):

there should be 2 solutions.

OpenStudy (anonymous):

x = -2

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

\[ x =- \frac{1}{3} \] x = -2

OpenStudy (anonymous):

sorry had a brain fart :-D

OpenStudy (anonymous):

* Bhaskara: x = ( -b(+-)√(b^2-4ac) )/2a * Your equation: 3x^2 + 7x + 2 = 0 * a = 3; b = 7; c = 2 * x = ( -7 (+-) √(7^2-4*3*2) ) / 2*3 * x = (-7 (+-) √25) ) / 6 * x = (-7 (+-) 5 ) / 6 * x1 = (-7+5)/6 = -1/3 * x2 = (-7-5)/6 = -2

OpenStudy (anonymous):

oh Ive only heard it called the quadratic formula.

OpenStudy (anonymous):

More about Bhaskara: http://en.wikipedia.org/wiki/Bh%C4%81skara_II

OpenStudy (anonymous):

I like hear more about people than formulas! ;)